You are considering buying a company for $325,000. If you expect the business to earn $113,000 per year, how long is the discounted payback period if your MARR is 10%? (in years, round to one decimal place) Please answer it by hand not using excel
Answer #1

Discounted payback period is time(years) to recover initial cost by periodic future cash flow in present value terms.

Provided,

Initial Cost (I) = $325,000

Annual Cash Inflow (A) = $113,000

MARR(i) = 10%

Now, we compute the cumulative present value of annual cash flows till it crossed the initial cost.

Year Cash Flow Present value of Cash flow Cum. Cash Flow
0 -$325,000 -$325,000 -$325,000.00
1 $113,000 113. 000 (1 + 0.1)1 = $102, 727.27 (-$325,000.00+$102,727.27) = -$222,272.73
2 $113,000 113.000 1400g = $593 388 47 (-$222,272.73+$93,388.43)= -$128,884.28
3 $113,000 113.000 (1+013 = $84.898.57 (-$128,884.28+$84,898.57)= -$43,985.73
4 $113,000 113,000 (1 + 0.1)4 = $77, 180,52 (-$43,985.73+$77,180.52)= $33194.80

With above calculation, we can say Initial cost is fully recovered between year 3 and Year 4.

Thus,

Discounted payback period:

-Cum. PV cash flow year 3 = 3+ - PV Cash flow in year 4

=3+- --43.985.73) 77.180.52

=3+ 43.985.73 77.180.52

=3+0.57

= 3.57 years

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.